OpenStudy (anonymous):

∫_0^π▒∫_y^π▒〖sinx/x dx dy〗

OpenStudy (anonymous):

You have: \[\int^{\pi}_{0} \int^{\pi}_{y} \frac{\sin(x)}{x} dx dy\] Obviously you realize this is tough to integrate. So you need to draw your region (your limits of integration, and reorder the integral do dydx. That MAY make the integral easier to solve. Try that and tell me what you get :P

OpenStudy (zarkon):

once you change the limits of integration it is really easy (you wont be able to do it the way it is currently written) you should get 2

OpenStudy (anonymous):

Is this cosine integral?